OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5).

OpenStudy (amistre64):

these do get rather repetitive:) same formula, same concept. y = m(x-x0)+y0 ; m=3, x0=2, and y0=5

OpenStudy (amistre64):

y = 3(x-2)+5 y = 3x -6+5 y = 3x-1

OpenStudy (amistre64):

i spose i could spice it up and say: \[y = mx -mx0 +y0\] :)

OpenStudy (anonymous):

... Stop doing them for her and they'll stop being repetitive.

OpenStudy (anonymous):

This is what I meant about the medals. She gives you a medal and you're back on the leash.

OpenStudy (amistre64):

it aint the medals; its the attitude. And yes, she prolly could still care less about how the answer is received; and I could still care less about any medals. But the illusion of being civil is what really count lol

OpenStudy (anonymous):

I don't think so. I think you can still encourage her to learn them on her own while being civil.

OpenStudy (anonymous):

And that's (IMO) the important part.

OpenStudy (amistre64):

:) its a good opinion to have I might say